OpenStudy (anonymous):

Check my answer: Use mathematical induction to prove the statement is true for all positive integers n. 8 + 16 + 24 + ... + 8n = 4n(n + 1)

OpenStudy (anonymous):

8+16+24+...+8k=4k(k+1) When we start off with k, we get 8k. Sk would equal 4k(k+1) When we use the next integer, k+1, we get 8(k+1). S(k+1) would equal 8k+8(k+1)=8k+8k+8=16k+8=8(2k+1)

OpenStudy (anonymous):

Then I'm stuck there.

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

let p(n)=8+16+24+....+8n=4n(n+1) for n=1,p(1)=4*1(1+1)=4*2=8 hence p(1) is true. assume that p(k) is true. 8+16+24+...+8k=4k(k+1) adding both sides 8(k+1),we get 8+16+24+...+8k+8(k+1)=4k(k+1)+8(k+1)=4(k+1)(k+2)=p(k+1)(k+1+1)=p(k+1) therefore if p(k) is true then p(k+1) is also true. Hence by induction p(n) is true for all positive integers.